OpenStudy (anonymous):

Evaluate the indefinite integral: 4xsin(x)dx

OpenStudy (anonymous):

i believe the easiest way is to use by parts

OpenStudy (anonymous):

aka tabulars method, are you in calculus 2?

OpenStudy (anonymous):

i got it down to -4xcos(x)+4cos(x) but i don't know if thats right

OpenStudy (anonymous):

yeah calculus 2

OpenStudy (anonymous):

i think the second should be a sin x

OpenStudy (anonymous):

-4xcos(x)+4sin(x)

OpenStudy (anonymous):

thats right

myininaya (myininaya):

\[\int\limits_{}^{}x \sin(x) dx=x \cdot( -\cos(x))-\int\limits_{}^{} 1 \cdot (-\cos(x)) dx=-x \cos(x)+\int\limits_{}^{}\cos(x) dx\] \[=-x \cos(x)+\sin(x)+C\] => \[4 \int\limits_{}^{} x \sin(x) dx=-4x \cos(x)+4 \sin(x)+K\]

OpenStudy (anonymous):

I''ll show you tabulars in a drawing